The measure of the supplement of an angle is 4 times the measure of its complement. Find the measure of one half its complement. I think it's 60 but i need explanation plzz. Thank you
Answer by oberobic(2304)

a = angle
180 -a = supplement
90 -a = complement
.
supplement = 4*complement
.
180-a = 4*(90-a)
180-a = 360 -4a
Add 4a to both sides
3a +180 = 360
Subtract 180 from both sides
3a = 180
Divide both sides by 3
a = 60.
.
Done.